please help me solve this word problem, Find three consecutive integers such that twice the first minus the third is 3 less than the second.
Answer by stanbon(75887)

Find three consecutive integers such that twice the first minus the third is 3 less than the second.
------
1st: x-1
2nd: x
3rd: x+1
-----------------
Equation:
2(x-1)-(x+1) = x-3
------
2x-2-x-1 = x-3
------
x -3 = x-3
-----
Ans: True for all values of "x"::
Example:: 2(7)-9 = 8-3
14-9 = 5
5 = 5
-----
Cheers,
Stan H.
------------